# score-based generative model SNR signal-to-noise ratio

### v2

@@ III-A Multi-step streaming diffusion
One approach to circumvent this is the Diffusion Buffer [lay2025diffusion], see Section II-D. In a follow-up preprint [lay2025diffusionjpre], the authors show that with a different training loss and direct-prediction inference, frames can be output earlier to reduce latency. However, this inference approximates the reverse process that the DNN was trained for, with the quality of the approximation decreasing as one decreases the latency. This results in significant output quality degradations when targeting the same algorithmic latency as our method, see Table II ($d=0$), suggesting the need for a different approach in low-latency settings.
@@ III-B Efficient real-time streaming diffusion inference
where $\star$ is the (possibly dilated) convolution evaluated only for a single output frame, yielding an output in $\mathbb{R}^{c_{o,l}\times 1}$, and each $\varphi_{l}$ represents arbitrary extra operations between convolutions assumed to be point-wise in time. We can then set $\tilde{Y}_{L}[t]$ as the output frame of this single DNN call. Note that all $k_{l},d_{l}$ can be chosen arbitrarily and independently without affecting the efficiency of this scheme. This idea has also been explored for real-time video processing [hedegaard2022cins]. The authors note that strided convolutions incur a delay to their respective downstream module and hence incur algorithmic latency, so we avoid their use here and set all time-wise convolution strides to 1.
@@
This scheme now performs the same computations as an offline implementation, only spread across physical time. Importantly, it yields exactly the same result—up to tiny numerical floating-point errors—as feeding the entire sequence to the causal model at once in an offline, batch-wise fashion. This fact enables the use of standard batched training and the subsequent approximation-free use of the trained weights for streaming inference, in contrast to, e.g., the Diffusion Buffer when outputting frames early for lower latency [lay2025diffusionjpre].
For future research, we note that this scheme can also be used with SDEs with one minor modification. For SDE solvers, a new noise sample $\varepsilon_{n}$ is drawn at every solver step $n=1,\ldots,N$. By redefining
$Y_{n}[t]:=\tilde{Y}_{n,L}[t]+\tilde{\sigma}_{n}\varepsilon_{n}[t]$ (16)
where $\tilde{\sigma}_{n}$ is the level of noise added back in according to the SDE solver and diffusion schedule, and treating this as another point-wise operation in time, no other changes to the inference are required. However, we only investigate ODE solvers in the following since we work with FM models.
@@ III-C DNN architecture
We design a custom frame-causal CNN without strided convolutions, as a modified variant of the NCSN++ architecture [song2021scorebased]. NCSN++ is a non-causal 2D U-Net architecture with multiple residual blocks per level, a progressive down/upsampling path, as well as non-causal attention layers. [richter2023speech] provides a detailed description of the specific NCSN++ configuration often used for SE. As in [richter2023speech], our DNN receives the audio signal in the form of a complex-valued STFT (a 2D time-frequency signal), with the real and imaginary parts mapped to two real-valued DNN input/output channels. We perform the following modifications and simplifications:
@@
    We keep the FIR filters for anti-aliased down- and upsampling [song2021scorebased] along frequency, but remove the FIR filtering along time.
@@
    We remove all attention layers. It is possible to include causal attention layers, but we do not follow this here for simplicity.
@@
    We replace the original GroupNorm [wu2018group] with a custom sub-band grouped BatchNorm (SGBatchNorm), inspired by [chang2021subspectral]. SGBatchNorm performs joint grouping along channels and frequencies and normalizes each group. We use four frequency groups, and the channel grouping of NCSN++ [song2021scorebased]. Prior work [richter2024causal] used time-cumulative group normalization, which determines statistics frame-by-frame as a time-varying IIR filter. In contrast, SGBatchNorm determines running-average batch statistics during training which are frozen and thus time-invariant during inference.
@@
    In contrast to [richter2024causal], we keep the progressive input down/upsampling paths [song2021scorebased], by simply using causal convolutions and the SGBatchNorm described above.
@@
    To fuse skip connections, we always employ addition, while NCSN++ [song2021scorebased] used addition on the downward and channel concatenation on the upward path. Our proposed addition significantly reduces the number of features on the upward path.
@@ III-D Predictive-generative speech enhancement
Inspired by Lemercier et al. [lemercier2023storm], for our SE task we use a joint predictive-generative approach, where an initial predictor network $D_{\eta}$ with parameters $\eta$ is first trained to map $Y$ to an initial estimate $Z:=D_{\phi}(y)$, and $Z$ is then used instead of $Y$ for defining and training the generative model. For $D_{\eta}$, we use the same DNN architecture as for the FM model, but remove all diffusion time conditioning layers. To train $D_{\eta}$, we use the following loss:
@@
where $z:=\operatorname{iSTFT}(Z)$ and $\mathcal{L}_{\text{MR-STFT}}$ is a multi-resolution magnitude STFT $L_{1}$ loss similar to [yamamoto2020parallelwavegan], using a set of $\operatorname{STFT}_{w}$ with different windows $w$ with $N_{w}$ window configurations. We use $N_{w}=4$ Hann windows with $W_{\text{ana}}\in\{256,512,768,1024\}$ and 50% overlap.
@@ III-E Custom learned low-NFE ODE solvers
Given a fixed budget of DNN evaluations per frame (NFE), we develop specialized learned Runge-Kutta ODE solvers (see Section II-C) to optimize the achievable quality without model retraining or finetuning, at virtually no increase of computations during inference. We train the scheme’s parameters $\{\mathbf{A},\mathbf{b},\mathbf{c}\}$, given a pretrained flow matching model $v_{\theta}$ with frozen weights $\theta$, by solving the ODE (1) with the current Runge-Kutta (RK) scheme and treating the final output as the clean speech estimate via $\hat{s}:=\operatorname{iSTFT}(X_{1})$. To optimize $\{\mathbf{A},\mathbf{b},\mathbf{c}\}$, we backpropagate through the whole solved ODE path and multiple DNN calls to determine the gradients, using the following loss:
$\mathcal{L}_{\text{RK}}$ $:=-\text{SpeechBERTScore}(\hat{s},s)$ (19)
$\quad+0.001\cdot\text{MRLogSpecMSE}(\hat{s},s)$ (20)
@@
where the motivation for using negative SpeechBERTScore [saeki24speechbertscore] is to reduce phonetic hallucinations/confusions that generative methods can suffer from [richter2024causal, deoliveira2025gibberish]. MRLogSpecMSE refers to a multi-resolution log-magnitude STFT mean squared error (MSE) loss similar to [yamamoto2020parallelwavegan] which we use to increase fine high-frequency detail, which FM models tend to lose in low-NFE settings [welker2025flowdec], see the log-spectral distance (LSD) metric in Table II. For this loss, we use window sizes $W_{\text{ana}}\in\{320,512,640\}$ and 75% overlap.
@@
We ensure by construction of the optimized parameters that $\sum_{j}a_{ij}=c_{i},\sum_{i}b_{i}=1$, and $0.05\leq b_{i}\leq 1$ for all $1\leq i,j\leq r$ to ensure that all model calls contribute to the final estimate. We clip $c_{i}\leq 0.85$ to avoid evaluating the FM model close to $\tau\approx 1$ where the training target (4) is unstable, and put a quadratic penalty loss on each $a_{ij}$ outside the range [2,2][-2,2]. We train $\{\mathbf{A},\mathbf{b},\mathbf{c}\}$ with the Adam optimizer at a learning rate of $10^{-3}$, a batch size of 10, and 2-second audio snippets, for 25,000 steps on our training dataset. For the 4 NFE available under our runtime budget in the SE task, we initialize $\{\mathbf{A},\mathbf{b},\mathbf{c}\}$ from Kutta’s four-stage 3/8 scheme [hairer1993solving]. For all other tasks, we have five NFE available due to the lack of a predictive DNN, and we use one Ralston-2 step followed by one Ralston-3 step [ralston1962runge] for initialization. Our learned schemes do not guarantee a convergence order since their coefficients do not necessarily follow the required algebraic conditions [hairer1993solving, Eq. (2.21)], which we argue is acceptable since our solvers do not need to be general-purpose. We list all RK parameters learned for each task in the supplementary material.
@@ III-F Model compression through weight decoupling
We make use of DNN compression methods to reduce the computational costs of each network call. Using this, we aim to improve the quality given a fixed computational budget, or to decrease the number of computations at some reasonable decrease in output quality. There are various methods to this end including weight pruning, quantization, or model distillation, but here we specifically follow the decoupling approach of Guo et al. [guo2018decoupling] to approximate the large 2D convolution weight tensors $\mathbf{W}\in\mathbb{R}^{n_{o}\times n_{i}\times k_{h}\times k_{w}}$ with output/input channels $n_{o}$, $n_{i}$ and kernel size $k_{h}\times k_{w}$, since such convolutions incur most of the computational effort in our networks.
@@
We apply this weight compression to all $3\times 3$ Conv2d layers with $n_{o}\geq 9$ in our network. As also proposed in [guo2018decoupling] we then fine-tune the compressed models for 25,000 training steps. We deviate slightly from the authors’ representation, choosing $U\sqrt{S}$ for the depthwise and $\sqrt{S}V^{\top}$ for the pointwise weights to spread the singular values across the two layers and improve fine-tuning stability.
@@ IV-A1 Speech Enhancement
For our speech enhancement (SE) task, the signal corruption model is $y=s+n$, where $s$ is the clean audio and $n$ is some uncorrelated background noise. As the dataset, we use EARS-WHAM v2333see https://github.com/sp-uhh/ears_benchmark for the v2 release. [richter2024ears], downsampled to 16 kHz. We also use the EARS-WHAM v2 clean utterances as the dataset for all following tasks except dereverberation.
@@ IV-A2 Dereverberation
We investigate speech dereverberation using the EARS-Reverb v2 dataset [richter2024ears]. The signal corruption model is y=shy=s*h, where * indicates time-domain convolution and $h\in\mathbb{R}^{l_{h}}$ is a sampled room impulse response (RIR).
@@ IV-A3 Codec Post-Filtering
Inspired by ScoreDec [wu2024scoredec] and FlowDec [welker2025flowdec], we investigate the use of Stream.FM as a post-filter for a low-bitrate speech codec. FlowDec [welker2025flowdec] introduces FM-based generative post-filtering, proposing a non-adversarially trained variant of the neural codec DAC [kumar2023dac]. Since DAC is non-causal and computationally expensive, we use the Lyra V2 codec [^4] https://github.com/google/lyra instead, which is built for streaming speech coding on consumer devices. Lyra V2 produces one frame every 20 ms at our chosen bitrate of 3.2 kbit/s. To align the codec frames with our model, we change the STFT parameters to use 40 ms windows and 20 ms hops.
@@ IV-A5 STFT Phase Retrieval
Peer et al. have shown with DiffPhase [peer2023diffphase] that SGMSE+ [richter2023speech] can be modified to solve an STFT PR task, resulting in very high reconstruction quality. We extend this idea to our streaming setting, using only 50% STFT overlap instead of the 75% overlap used in DiffPhase. We compare our method against a family of streaming STFT PR algorithms proposed by Peer et al. [peer2024flexible].
@@ IV-A6 Mel Vocoding
As shown in [welker2025realtime], the ideas of DiffPhase [peer2023diffphase] can be extended to streaming Mel vocoding through a small change in the corruption model. Instead of treating the phaseless STFT magnitudes $|X|$ as the corrupted signal [peer2023diffphase], we additionally subject them to a lossy Mel compression as follows:
@@
where $M\in\mathbb{R}^{F_{\text{mel}}\times F_{\text{STFT}}}$ is the Mel matrix mapping STFT frames to Mel frames, $M^{\dagger}$ is its Moore-Penrose pseudoinverse, and $X[t]$ denotes the single magnitude spectrogram frame at frame index $t$. We follow the Mel configuration of HiFi-GAN [kong2020hifigan] in the 16 kHz variant from SpeechBrain [ravanelli2024speechbrain], but to reduce the latency, we use 32 ms windows instead of 64 ms while keeping the 16 ms hop length.
@@ IV-B Data and data representation
We use the EARS dataset [richter2024ears] as the basis of all our problem variants and model trainings, resampled to a sampling frequency of $f_{s}=$ 16 kHz. Unless otherwise noted, we use an STFT with a 512-point periodic $\sqrt{\text{Hann}}$-window (32 ms), a 256-point hop length (16 ms, 50% overlap), and magnitude compression with exponent $\alpha=0.5$ as in [welker2022speech, richter2023speech]. Different from these works, we use an orthonormal STFT and do not apply an additional scaling. Since a 512-point window leads to 257 frequency bins, for ease of DNN processing, we discard the Nyquist band to retrieve frames with 256 frequency bins, and pad it back with zeros before applying the inverse STFT.
For SE, similar to [richter2024causal], during training we peak-normalize $s$ and $y$ independently and apply a random negative gain between -12 and 0 dB to $y$, so that the model learns to perform automatic gain control. For all other tasks, we normalize $s$ and $y$ jointly based on the peak magnitude of $y$, assuming that a roughly constant input-output level relationship is available in these tasks. As the FM process hyperparameter $\Sigma_{y}$ (3), we empirically set scalar $\sigma_{y}=0.05$ for SE, $\sigma_{y}=0.25$ for STFT PR and Mel vocoding, and $\sigma_{y}=0.35$ for dereverberation and codec post-filtering. For BWE, we follow [welker2025flowdec] and determine a heuristic per-frequency-band diagonal covariance matrix $\Sigma_{y}$ to avoid adding noise in the preserved low-frequency bands and to allow easier regeneration of the low-energy high-frequency bands. We set $\sigma_{\min}=0.001$ for all tasks except BWE where $\Sigma_{\min}=0.001\cdot\Sigma_{y}$.

### v3

@@ III-A Multi-step streaming diffusion
One approach to circumvent this is the Diffusion Buffer [27], see Section I-A. In a follow-up preprint [28], the authors show that with a different training loss and direct-prediction inference, frames can be output earlier to reduce latency. However, this approximates the reverse process that the DNN was trained for, with the approximation quality decreasing as one decreases the latency. This results in significant quality degradations when targeting the same algorithmic latency as our method, see Table II ($d=0$), suggesting the need for a different approach in low-latency settings.
@@ III-B Efficient real-time streaming diffusion inference
where $\star$ is the (possibly dilated) convolution evaluated only for a single output frame, yielding an output in $\mathbb{R}^{c_{o,l}\times 1}$, and each $\varphi_{l}$ represents arbitrary extra operations between convolutions assumed to be point-wise in time. We can then set $\tilde{Y}_{L}[t]$ as the output frame of this single DNN call. Note that all $k_{l},d_{l}$ can be chosen arbitrarily and independently without affecting the efficiency of this scheme. This idea has also been explored for real-time video processing [16]. The authors note that strided convolutions incur a delay to their respective downstream module and hence incur algorithmic latency, so we avoid their use here and set all time-wise convolution strides to 1.
@@
This scheme now performs the same computations as an offline implementation, only spread across physical time. Importantly, it yields exactly the same result—up to tiny numerical floating-point errors—as feeding the entire sequence to the causal model at once in an offline, batch-wise fashion. This fact enables the use of standard batched training and the subsequent approximation-free use of the trained weights for streaming inference, in contrast to, e.g., the Diffusion Buffer when outputting frames early for lower latency [28].
We note that this scheme can also be used with SDEs with one minor modification. For SDE solvers, a new noise sample $\varepsilon_{n}$ is drawn at every solver step $n=1,\ldots,N$. By redefining $Y_{n}[t]:=\tilde{Y}_{n,L}[t]+\tilde{\sigma}_{n}\varepsilon_{n}[t]$, where $\tilde{\sigma}_{n}$ is the level of noise added according to the SDE solver and diffusion schedule, and treating this as another point-wise operation in time, no other changes to the inference are required.
@@ III-C DNN architecture
We design a custom frame-causal CNN without strided convolutions, as a modified variant of the NCSN++ architecture [54]. NCSN++ is a non-causal 2D U-Net architecture with multiple residual blocks per level, a progressive down/upsampling path, as well as non-causal attention layers. For a detailed description and illustration of the specific NCSN++ configuration we build upon, we refer the reader to SGMSE+ [47]. As in [47], our DNN receives the audio signal as a complex-valued STFT (a 2D time-frequency signal), with the real and imaginary parts mapped to two real-valued DNN input/output channels. We perform the following modifications and simplifications:
@@
    We keep the FIR filters for anti-aliased down- and upsampling [54] along frequency, but remove the FIR filtering along time.
@@
    We remove all attention layers. It is in principle possible to include causal attention layers while staying streaming-capable by using causal attention with a fixed-size window for constant runtime and memory complexity, but for simplicity, our main models do not use attention. We briefly investigate adding such a layer in our ablation study, see Section V-I.
@@
    We replace the original GroupNorm [65] with a custom sub-band grouped BatchNorm (SGBatchNorm), inspired by [6]. SGBatchNorm performs joint grouping along channels and frequencies and normalizes each group. We use four frequency groups, and the channel grouping of NCSN++ [54]. Prior work [48] used time-cumulative group normalization, which determines statistics frame-by-frame as a time-varying IIR filter. In contrast, SGBatchNorm determines running-average batch statistics during training which are frozen and thus time-invariant during inference.
@@
    In contrast to [48], we keep the progressive input down/upsampling paths [54], by simply using causal convolutions and the SGBatchNorm described above.
@@
    To fuse skip connections, we always employ addition, while NCSN++ [54] used addition on the downward and channel concatenation on the upward path. Our proposed addition significantly reduces the number of features on the upward path.
@@
All modifications described above are implemented in the DNN architecture we make available in our public codebase. Further details on engineering-level optimizations of our models towards real-time inference can be found in Section S.VIII of the supplementary material.
@@ III-D Predictive-generative speech enhancement
Inspired by Lemercier et al. [31], for our SE task we use a joint predictive-generative approach, where an initial predictor network $D_{\eta}$ with parameters $\eta$ is first trained to map $Y$ to an initial estimate $Z:=D_{\phi}(y)$, and $Z$ is then used instead of $Y$ for defining and training the generative model. For $D_{\eta}$, we use the same DNN architecture as for the FM model, but remove all diffusion time conditioning layers. To train $D_{\eta}$, we use the following loss:
@@
where $z:=\operatorname{iSTFT}(Z)$ and $\mathcal{L}_{\text{MR-STFT}}$ is a multi-resolution magnitude STFT $L_{1}$ loss similar to [66], using a set of $\operatorname{STFT}_{w}$ with different windows $w$ with $N_{w}$ window configurations. We use $N_{w}=4$ Hann windows with $W_{\text{ana}}\in\{256,512,768,1024\}$ and 50% overlap.
@@ III-E Custom learned low-NFE ODE solvers
Given a fixed budget of DNN evaluations per frame (NFE), we develop specialized Learned Runge-Kutta (LRK) ODE solvers (see Section II-C) to optimize the achievable quality without model retraining or finetuning, at virtually no increase of computations during inference. Conceptually, our proposal is related to [14] from numerical literature, but differs in the following ways: we use loss functions specific to speech processing instead of a simple mean squared error (MSE); we set $\Delta\tau=1$ to guarantee a fixed low NFE; since intermediate results are not valid clean speech estimates, we only calculate the loss on the single endpoint $\hat{x}_{1}=\hat{s}$; our integrated function $v_{\theta}$ is high-dimensional and given by a neural network, rather than a simple function from low-dimensional numerical ODE problems. We train the scheme’s parameters $\{\mathbf{A},\mathbf{b},\mathbf{c}\}$, given a pretrained flow matching model $v_{\theta}$ with frozen weights $\theta$, by solving the ODE (1) with the current Runge-Kutta (RK) scheme and treating the final output as the clean speech estimate via $\hat{s}:=\operatorname{iSTFT}(X_{1})$. To optimize $\{\mathbf{A},\mathbf{b},\mathbf{c}\}$, we backpropagate through the whole solved ODE path and multiple DNN calls to determine the gradients, using the following loss:
[…]
@@
where SBS refers to SpeechBERTScore [51] and MR-LS-MSE refers to a multi-resolution log-magnitude STFT MSE loss similar to [66]. The motivation for using negative SpeechBERTScore [51] is to reduce phonetic hallucinations that generative methods can suffer from [48, 8], and the motivation for MR-LS-MSE is to increase fine high-frequency detail, which FM models tend to lose in low-NFE settings [60]. For this loss, we use window sizes $W_{\text{ana}}\in\{320,512,640\}$ and 75% overlap. We further consider a loss oriented towards signal fidelity, replacing the SpeechBERTScore term with a differentiable [^4] https://github.com/audiolabs/torch-pesq PESQ loss [38]:
[…]
@@
We ensure by construction of the optimized parameters that $\sum_{j}a_{ij}=c_{i},\sum_{i}b_{i}=1$, and $0.05\leq b_{i}\leq 1$ for all $1\leq i,j\leq r$ to ensure that all model calls contribute to the final estimate. The constraint $\sum_{i}b_{i}=1$ guarantees the consistency of our schemes, as well as convergence with order of (at least) 1 [5]. The region of absolute stability of each LRK scheme can be found in Fig. S5 of the supplementary material. Our LRK schemes do not guarantee a convergence order larger than 1 since the learning-based procedure of determining the coefficients makes it unlikely that they will satisfy the required algebraic conditions [15, Eq. (2.21)] for order-2 and above, which we argue is acceptable since our solvers do not need to be general-purpose. We find that they nonetheless transfer reasonably well, see Section V-G. We clip $c_{i}\leq 0.85$ to avoid evaluating the FM model close to $\tau\approx 1$ where the training target (4) is unstable, and put a quadratic penalty loss on each $a_{ij}$ outside the range [2,2][-2,2]. We train $\{\mathbf{A},\mathbf{b},\mathbf{c}\}$ with the Adam optimizer at a learning rate of $10^{-3}$, a batch size of 10, and 2-second audio snippets, for 25,000 steps on our training dataset. For the 4 NFE available under our runtime budget in the SE task, we initialize $\{\mathbf{A},\mathbf{b},\mathbf{c}\}$ from Kutta’s four-stage 3/8 scheme [15]. For all other tasks, we have five NFE available due to the lack of a predictive DNN, and we use one Ralston-2 step followed by one Ralston-3 step [44] for initialization. We list the RK parameters learned for each task in Section S.IX-B of the supplementary material.
@@ III-F Model compression through weight decoupling
We make use of DNN compression methods to reduce the computational costs of each network call. Using this, we aim to improve the quality given a fixed computational budget, or to decrease the number of computations at some reasonable decrease in output quality. There are various methods to this end including weight pruning, quantization, or model distillation, but here we specifically follow the decoupling approach of Guo et al. [13] to approximate the large 2D convolution weight tensors $\mathbf{W}\in\mathbb{R}^{n_{o}\times n_{i}\times k_{h}\times k_{w}}$ with output/input channels $n_{o}$, $n_{i}$ and kernel size $k_{h}\times k_{w}$, since such convolutions incur most of the computational effort in our networks.
@@
We apply this weight compression to all $3\times 3$ Conv2d layers with $n_{o}\geq 9$ in our network. As also proposed in [13] we then fine-tune the compressed models for 25,000 training steps. We deviate slightly from the authors’ representation, choosing $U\sqrt{S}$ for the depthwise and $\sqrt{S}V^{\top}$ for the pointwise weights to spread the singular values across the two layers and improve fine-tuning stability.
@@ IV-A1 Speech Enhancement
For our speech enhancement (SE) task, the signal corruption model is $y=s+n$, where $s$ is the clean audio and $n$ is some uncorrelated background noise. As the dataset, we use EARS-WHAM v2555see https://github.com/sp-uhh/ears_benchmark for the v2 release. (EWv2) [49], downsampled to 16 kHz. We also use the EARS-WHAM v2 (EWv2) clean utterances as the dataset for all following tasks except dereverberation.
@@ IV-A2 Dereverberation
We investigate speech dereverberation using the EARS-Reverb v2 dataset [49]. The signal corruption model is y=shy=s*h, where * indicates time-domain convolution and $h\in\mathbb{R}^{l_{h}}$ is a sampled room impulse response (RIR).
@@ IV-A3 Codec Post-Filtering
Inspired by ScoreDec [64] and FlowDec [60], we investigate the use of Stream.FM as a post-filter for a low-bitrate speech codec. FlowDec [60] introduces FM-based generative post-filtering, proposing a non-adversarially trained variant of the neural codec DAC [26]. Since DAC is non-causal and computationally expensive, we use the Lyra V2 codec [^6] https://github.com/google/lyra instead, which is built for streaming speech coding on consumer devices. Lyra V2 produces one frame every 20 ms at our chosen bitrate of 3.2 kbit/s. To align the codec frames with our model, we change the STFT parameters to use 40 ms windows and 20 ms hops.
@@ IV-A5 STFT Phase Retrieval
Peer et al. have shown with DiffPhase [40] that SGMSE+ [47] can be modified to solve an STFT PR task, resulting in very high reconstruction quality. We extend this idea to our streaming setting, using only 50% STFT overlap instead of the 75% overlap used in DiffPhase. We compare our method against the non-causal DiffPhase [40], which we reconfigure with $W_{\text{syn}}=510,H_{\text{syn}}=255$ for exactly 50% overlap. We use these STFT parameters for compatibility with the DiffPhase code and DNN. We retrain DiffPhase using the same data, batchsize, and number of optimizer steps as for our method. We further compare against a family of streaming STFT PR algorithms proposed by Peer et al. [41].
@@ IV-A6 Mel Vocoding
As we show in [61], the ideas of DiffPhase [40] can be extended to streaming Mel vocoding through a small change in the corruption model. Instead of treating the phaseless STFT magnitudes $|X|$ as the corrupted signal [40], we additionally subject them to a lossy Mel compression as follows:
@@
where $M\in\mathbb{R}^{F_{\text{mel}}\times F_{\text{STFT}}}$ is the Mel matrix mapping STFT frames to Mel frames, $M^{\dagger}$ is its Moore-Penrose pseudoinverse, and $X[t]$ denotes the single magnitude spectrogram frame at frame index $t$. We follow the Mel configuration of HiFi-GAN [22] in the 16 kHz variant from SpeechBrain [45], but to reduce the latency, we use 32 ms windows instead of 64 ms while keeping the 16 ms hop length.
@@ IV-B Data and data representation
We use the EARS dataset [49] as the basis of all our problem variants and model trainings, resampled to a sampling frequency of $f_{s}=$ 16 kHz. In evaluations for SE, we also use the VoiceBank-DEMAND (VB-DMD) [56] dataset, also resampled to 16 kHz. Unless otherwise noted, we use an STFT with a 512-point periodic $\sqrt{\text{Hann}}$-window (32 ms), a 256-point hop length (16 ms, 50% overlap), and magnitude compression with exponent $\alpha=0.5$ as in [62, 47]. Different from these works, we use an orthonormal STFT and do not apply an additional scaling. Since a 512-point window leads to 257 frequency bins, for ease of DNN processing, we discard the Nyquist band to retrieve frames with 256 frequency bins, and pad it back with zeros before applying the inverse STFT.
For SE, similar to [48], during training we peak-normalize $s$ and $y$ independently and apply a random negative gain between -12 and 0 dB to $y$, so that the model learns to perform automatic gain control. For all other tasks, we normalize $s$ and $y$ jointly based on the peak magnitude of $y$, assuming that a roughly constant input-output level relationship is available in these tasks. As the FM process hyperparameter $\Sigma_{y}$ (3), we empirically set scalar $\sigma_{y}=0.05$ for SE, $\sigma_{y}=0.25$ for STFT PR and Mel vocoding, and $\sigma_{y}=0.35$ for dereverberation and codec post-filtering. For BWE, we follow [60] and determine a heuristic per-frequency-band diagonal covariance matrix $\Sigma_{y}$ to avoid adding noise in the preserved low-frequency bands and to allow easier regeneration of the low-energy high-frequency bands. We set $\sigma_{\min}=0.001$ for all tasks except BWE where $\Sigma_{\min}=0.001\cdot\Sigma_{y}$.